can you please help me with this question. I have been confused for about three hours and I would really appreciate it if you would help me. thanks....
x^2 - 7x - 18 = 0

Answer by checkley71(8403)

X^2-7X-18=0
(X-9)(X+2)=0
X-9=0
X=9 ANSWER.
X+2=0
X=-2 ANSWER.